# Loss Landscape Degeneracy Drives Stagewise Development in Transformers

### v2

@@ C.2.4 Prefix matching score
The prefix matching score from Olsson et al. (2022) is defined similarly to the previous-token matching score. Given a sequence $[A][B]\ldots[A]$, the prefix matching score of a particular attention head is how much the attention head attends back to the first instance of $[A]$ when at the second instance of $[A]$.
We compute this score using a synthetic data-generating process. We first generate 10k fixed random sequences of length 128. The first token is always a special “beginning of string" token and the $[A]$ and $[B]$ tokens are selected and placed randomly. One $[A]$ token is placed in the first half of the sequence, the other is placed in the second half, and the $[B]$ token is placed directly after the first $[A]$ token. The remaining tokens are randomly sampled from the tokenizer vocabulary, excluding the $[A]$, $[B]$, and beginning of string tokens.
For each sample in this synthetic dataset, we measure the attention score that each attention head assigns to the earlier instance of $[A]$ from the latter instance of $[A]$. These scores are averaged across the dataset to produce the prefix matching score for that attention head at a given checkpoint. The progression of prefix matching scores over time can be seen in Figure 4(e).
@@ Appendix D Developmental analysis of regression transformers
In this section, we present further evidence on the behavioral (Section D.1) and structural (Section D.2) development of the transformer in the setting of in-context linear regression.
@@ D.1.1 Task prior score
Figure D.1 depicts (a) the mean square distance between the model’s predictions and the zero prediction in addition to (b) the mean square distance between the model’s predictions and the “task prior” prediction, using the component-wise averaged $\overline{\mathbf{t}}$ over the set of tasks encountered during training. For all models, the minimum distance to the task prior prediction is lower than the minimum distance to the zero prediction. Hence, we call stage LR1 “learning the task prior” rather than simply learning the zero prediction.
Figure D.1: Learning the task prior is universal across models trained on very different data distributions. Each line represents a model trained on a data distribution with a different number of $M$ distinct tasks (“task diversity” in Raventós et al., 2023). In addition to taking a finite $M$, the models depicted here differ from the other models considered in this paper in that the former were trained with a maximum learning rate of $0.01$, and the models (inadvertently) lack an output matrix after the multi-head attention layer.
@@ D.1.3 OOD generalization
Note that these inputs and tasks are not out-of-distribution in the sense of coming from a distribution with a different support than the training distribution. However, the samples drawn from these “extreme” distributions are exponentially suppressed by the original training distribution. Figure D.3 plots the normalized MSE for these two distributions over training time.
@@ D.2.1 Embedding
The embedding matrix $W_{E}$ is a linear transformation from $\mathbb{R}^{D+1}\to\mathbb{R}^{{d_{\text{embed}}}}$. Plotting the $D+1$ singular values of this matrix, we notice that the embedding partially loses one of its components starting at the end of LR2 (Figure D.4a).
The input “tokens” $x_{k}$ span a $D$-dimensional subspace of the $(D+1)$-dimensional “token space.” The target tokens $y_{k}$ span an orthogonal $1$-dimensional subspace. The collapse of one of the embedding matrix’s singular values means that the model learns to redundantly encode the inputs and targets in the same $D$-dimensional subspace of the space of residual stream activations. The almost order of magnitude separation in the magnitudes of the square singular value means that the $(D+1)$th component of the token embedding explains only 2.9% of the variance in activations of the residual stream immediately after the embedding, whereas the dominant components explain roughly 24% each.
@@ D.2.2 Positional encoding
The positional encoding goes through a similar collapse to the unembedding starting during the second part of LR2 and continuing into LR3 (Figure D.4b). Additionally, throughout these stages, the subspace spanned by the embedding becomes more aligned with the subspace spanned by the positional encoding (Figure D.4c).
@@
For the same reason as with the token embedding, a decrease in the dimensionality of the subspace occupied by activations reduces the effective number of dimensions and thus the learning coefficient. This occurs both as the positional encoding’s effective dimensionality decreases (vanishing singular values, Figure D.4b) and as the token embedding subspace and positional embedding subspace align (increasing cosine similarity, Figure D.4b).
@@ D.2.3 Attention collapse
Over the course of training, we observe that some attention heads learn to attend solely (soft attention becomes hard attention) and consistently to certain positions (the attention pattern becomes content-independent). We call this phenomenon attention collapse in parallel with the other observed forms of collapse. Not only does this potentially contribute to a decrease in the LLC, but it also makes the attention heads identifiable: we find a self-attention head, previous-attention heads, previous-$x$-attention heads, and previous-$y$-attention heads.
@@
To quantify attention hardness, we use the attention entropy score (Ghader and Monz, 2017; Vig and Belinkov, 2019). Given the attention pattern $\alpha_{k,k^{\prime}}^{(b,h)}$ for how much token $k$ in head $h$ in block $b$ attends back to token $k^{\prime}$, its attention entropy score $H_{k}^{(b,h)}$ is the Shannon entropy over preceding indices $k^{\prime}<k$,
@@
This accounts for the entropy being calculated over different numbers of tokens and is displayed in Figure D.5. Notably, the identified stages line up closely to stages of these attention entropy curves.
Figure D.5: Attention hardening as measured by the normalized attention entropy score (Section D.2.3). Block 1 heads 1y/3y and block 2 head 1y harden over training. In combination with the fact that these attention heads become less variable (Figure D.6), this may contribute to a decrease in the LLC (discussed in Section D.2.3) The x-components of the attention heads remain much softer over the entire training run.
@@
Accomplishing constant attention requires the presence of biases in the query and key transformations, or if there is no bias (as is the case for the models we investigated), requires attending to the positional embedding. With the Shortformer-style positional encoding used for the language models (Section F.1.1), this is straightforward: the positional information is injected directly into the key and weight matrices. With the in-context linear regression models, where the positional embedding is added to the residual stream activations, this is less straightforward: achieving constant attention requires separating residual stream activations into orthogonal positional- and input-dependent subspaces, then reading from the former with the query and key weight matrices.
@@
To quantify how constant the attention pattern is, we use measure attention variability (Vig and Belinkov, 2019),
@@
where the division by $2$ ensures the variability lies in the range $[0,1]$. This is displayed in Figure D.6. These reveal that though attention hardness and variability are independent axes of differentiation, empirically, we observe that hard attention is correlated with low variability.
Figure D.6: Attention variability over time. The heads that develop hard attention in Figure D.5 (block 1 heads 1y, 3y, and 4y) also become less variable over time.
@@
Previous-token attention is measured the same as in the language model setting (Section C.2) with one difference: we compute the previous-token score not over a synthetic dataset but over a validation batch.
@@
Several attention heads are easy to identify by virtue of being both concentrated and consistent. These are depicted in Figure D.7 and include: (B1H3y) previous-token heads (also present in the language model case), (B1H1y) previous-x, and (B1H4x, B2H1y) previous-y heads. Other training runs also include self-attention heads.
Figure D.7: Collection of attention heads identified by their consistent and recognizable attention patterns. Left to right: previous-$x$s head, previous-token head, previous-$y$s head, previous-$y$s head
@@
Suppose an attention head $h$ in block $b$ has the following constant attention pattern (after the softmax) $A^{(b,h)}=\sum_{i}\delta_{l(i)\,i}$. That is, for each token $i$, that attention head attends solely to a single earlier token $l(i)\leq i$ and no others. Restricting to single-head attention (the argument generalizes straightforwardly), the final contribution of this attention head to the residual stream is the following (Phuong and Hutter, 2022):
@@
Note that this argument does not hold for all hard and constant attention patterns. It holds solely for attention patterns that consistently ignore some earlier token across all positions, such as the previous-$x$ and previous-$y$ heads, but not the self-attention and previous-token heads. As discussed in Section D.2.6, it remains unclear what exactly the threshold for “ignoring” a token should be before it contributes to degeneracy and whether any of the heads we examine actually meet this threshold.
@@ D.2.4 Unembedding collapse
where we order the outputs so that the $y$-token corresponds to the 00th row. Because we are reading out a single $y$ component, we can express the unembedding transformation in terms of “effective" unembedding weights and biases
@@
In Figure D.8, we plot $(\gamma,\beta)$, $((W_{U})_{[0,:]},(b_{U})_{[0]})$, and $(\tilde{W}_{U},\tilde{b}_{U})$ as a function of training steps, along with the mean weight over time. These are 64- and 1-dimensional vectors, so we can display the entire set of components. During stage LR3 the majority of weights $\beta$ and $W_{U}$ “collapse” to zero. Additionally, the layer normalization biases temporarily experience a large increase in variance before returning to small values. Despite this, the mean of the linear weights, layer normalization biases, and effective weights remains remarkably constant and close to zero throughout the entire process.
@@ D.2.5 Layer normalization collapse
The “collapse” in layer normalization weights is not unique to the unembedding. As depicted in Figure D.9, this behavior occurs in all layer norms except for the second MLP. The biases also remain centered close to zero even as the variance in biases grows much larger. Unlike in the unembedding, these layers begin to change earlier (starting halfway through LR2).
@@ D.2.6 Degeneracy and development
In the previous subsections, we provide a set of theoretical arguments for how embedding collapse (Section D.2.1), layer normalization collapse (Section D.2.5), and attention collapse (Section D.2.3) can lead to an increase in degeneracy, even while leaving the implemented function unchanged.
@@ Appendix E One-layer language model experiments
We also trained and ran some experiments on a one-layer language model (see Section F.1.1 for details). We aggregate results for the one-layer language model here, mirroring the experiments for the two-layer language model where possible. The early development of the one-layer model has many parallels with the two-layer model. At a single stage boundary, just as it occurs in the two-layer model, the one-layer model minimizes its bigram score (see Section C.1.1), begins utilizing the positional embedding to noticeably improve performance (see Section C.2.1), and starts making sudden improvements to the same $n$-gram scores (see Section C.1.2). Remarkably this occurs at the same checkpoint as in the 2-layer model (at 900 training steps).
@@
Figure E.1: We train a one-layer transformer model in the language setting to compare with the two-layer model. The development of certain behavioral and structural metrics over time closely mirrors the development of the same metrics in the early stages of the two-layer language model. Top: test loss and LLC estimations over time for the one-layer attention-only transformer, compare with Figure 1(a). Bottom: bigram score, test loss with positional embedding ablated, and $n$-gram scores for the one-layer attention-only transformer, compare with Figure 4(a,b,c).
Figure E.2: A more detailed version of Figure E.1 for the one-layer language model. Top: Loss, LLC, and weight norm, along with an overlaid Gaussian process fit to these curves (red dotted lines). Bottom: Associated slopes, both numerically estimated finite differences (transparent blue) and of the Gaussian process (red dotted lined).
@@ F.1.1 Architecture
The language model architectures we consider are one- and two-layer attention-only transformers. They have a context length of 1024, a residual stream dimension of $d_{model}=256$, $H=8$ attention heads per layer, and include layer normalization layers. We also used a learnable Shortformer positional embedding (Press et al., 2021). The resulting models have a total of $d=3,091,336$ parameters for $L=1$ and $d=3,355,016$ parameters for $L=2$. We used an implementation provided by TransformerLens (Nanda and Bloom, 2022).
@@
Token Embedding Weights $1,280,000$
Positional Embedding Weights $262,144$
Layer 1 Layer Norm Weights $256$
Layer 1 Layer Norm Bias $256$
Layer 1 Attention Query Weights $65,536$
Layer 1 Attention Key Weights $65,536$
Layer 1 Attention Value Weights $65,536$
Layer 1 Attention Output Weights $65,536$
Layer 1 Attention Query Bias $256$
Layer 1 Attention Key Bias $256$
Layer 1 Attention Value Bias $256$
Layer 1 Attention Output Bias $256$
Layer 2 Layer Norm Weights N/A $256$
Layer 2 Layer Norm Bias N/A $256$
Layer 2 Attention Query Weights N/A $65,536$
Layer 2 Attention Key Weights N/A $65,536$
Layer 2 Attention Value Weights N/A $65,536$
Layer 2 Attention Output Weights N/A $65,536$
Layer 2 Attention Query Bias N/A $256$
Layer 2 Attention Key Bias N/A $256$
Layer 2 Attention Value Bias N/A $256$
Layer 2 Attention Output Bias N/A $256$
Final Layer Norm Weights $256$
Final Layer Norm Bias $256$
Unembedding Weights $1,280,000$
Unembedding Bias $5,000$
@@ F.1.2 Tokenization
For tokenization, we used a truncated variant of the GPT-2 tokenizer that cut the original vocabulary of 50,000 tokens down to 5,000 (Eldan and Li, 2023) to reduce the size of the model. We think this may contribute to the prominence of the the plateau at the end of LM1: the frequency of bigram statistics depends on your choice of tokens, and a larger tokenizer leads to bigrams that are individually much less frequent.
@@ F.1.3 Training
The models are trained on a single epoch over $50,000$ steps on $\sim$5 billion tokens using a resampled subset of the Pile (Gao et al., 2020; Xie et al., 2023) using a batch size of $100$. A snapshot was saved every $10$ steps for a total of $5000$ checkpoints, though a majority of analysis used checkpoints every 100 steps. The training time was around 6 GPU hours per model on an A100. Additional seeds were trained on v4 TPUs at around 1.5 TPU hours per model.
@@
$n$ Data # of training samples $5,000,000$
$T$ Data # of training steps $50,000$
@@
$L$ Model # of layers in the model $2$
@@
$\eta$ Optimizer Learning rate $0.001$
$\lambda_{\mathrm{wd}}$ Optimizer Weight Decay $0.05$
$\beta_{1,2}$ Optimizer Betas $(0.9,0.999)$

### v3

@@ C.2.4 Prefix matching score
[…]
@@ Appendix D Developmental analysis of regression transformers
In this section, we present further evidence on the behavioral (Section˜D.1) and structural (Section˜D.2) development of the transformer in the setting of in-context linear regression.
@@ D.1.1 Task prior score
Figure˜D.1 depicts (a) the mean square distance between the model’s predictions and the zero prediction in addition to (b) the mean square distance between the model’s predictions and the “task prior” prediction, using the component-wise averaged $\overline{\mathbf{t}missing}$ over the set of tasks encountered during training. For all models, the minimum distance to the task prior prediction is lower than the minimum distance to the zero prediction. Hence, we call stage LR1 “learning the task prior” rather than simply learning the zero prediction.
Figure D.1: Learning the task prior is universal across models trained on very different data distributions. Each line represents a model trained on a data distribution with a different number of $M$ distinct tasks (“task diversity” in Raventós et al., 2023). In addition to taking a finite $M$, the models depicted here differ from the other models considered in this paper in that the former were trained with a maximum learning rate of 0.010.010.01, and the models (inadvertently) lack an output matrix after the multi-head attention layer.
@@ D.1.3 OOD generalization
Note that these inputs and tasks are not out-of-distribution in the sense of coming from a distribution with a different support than the training distribution. However, the samples drawn from these “extreme” distributions are exponentially suppressed by the original training distribution. Figure˜D.3 plots the normalized MSE for these two distributions over training time.
@@ D.2.1 Embedding
The embedding matrix $W_{E}$ is a linear transformation from $\mathbb{R}^{D+1}\to\mathbb{R}^{{d_{\text{embed}}}}$. Plotting the $D+1$ singular values of this matrix, we notice that the embedding partially loses one of its components starting at the end of LR2 (Figure˜D.4a).
[…]
@@ D.2.2 Positional encoding
The positional encoding goes through a similar collapse to the unembedding starting during the second part of LR2 and continuing into LR3 (Figure˜D.4b). Additionally, throughout these stages, the subspace spanned by the embedding becomes more aligned with the subspace spanned by the positional encoding (Figure˜D.4c).
@@
For the same reason as with the token embedding, a decrease in the dimensionality of the subspace occupied by activations reduces the effective number of dimensions and thus the learning coefficient. This occurs both as the positional encoding’s effective dimensionality decreases (vanishing singular values, Figure˜D.4b) and as the token embedding subspace and positional embedding subspace align (increasing cosine similarity, Figure˜D.4b).
@@ D.2.3 Attention collapse
[…]
@@
To quantify attention hardness, we use the attention entropy score (Ghader & Monz, 2017; Vig & Belinkov, 2019). Given the attention pattern $\alpha_{k,k^{\prime}}^{(b,h)}$ for how much token $k$ in head $h$ in block $b$ attends back to token $k^{\prime}$, its attention entropy score $H_{k}^{(b,h)}$ is the Shannon entropy over preceding indices $k^{\prime}<k$,
@@
This accounts for the entropy being calculated over different numbers of tokens and is displayed in Figure˜D.5. Notably, the identified stages line up closely to stages of these attention entropy curves.
Figure D.5: Attention hardening as measured by the normalized attention entropy score (Section˜D.2.3). Block 1 heads 1y/3y and block 2 head 1y harden over training. In combination with the fact that these attention heads become less variable (Figure˜D.6), this may contribute to a decrease in the LLC (discussed in Section˜D.2.3) The x-components of the attention heads remain much softer over the entire training run.
@@
Accomplishing constant attention requires the presence of biases in the query and key transformations, or if there is no bias (as is the case for the models we investigated), requires attending to the positional embedding. With the Shortformer-style positional encoding used for the language models (Section˜F.1.1), this is straightforward: the positional information is injected directly into the key and weight matrices. With the in-context linear regression models, where the positional embedding is added to the residual stream activations, this is less straightforward: achieving constant attention requires separating residual stream activations into orthogonal positional- and input-dependent subspaces, then reading from the former with the query and key weight matrices.
@@
To quantify how constant the attention pattern is, we use measure attention variability (Vig & Belinkov, 2019),
@@
where the division by 222 ensures the variability lies in the range [0,1][0,1][ 0 , 1 ]. This is displayed in Figure˜D.6. These reveal that though attention hardness and variability are independent axes of differentiation, empirically, we observe that hard attention is correlated with low variability.
Figure D.6: Attention variability over time. The heads that develop hard attention in Figure˜D.5 (block 1 heads 1y, 3y, and 4y) also become less variable over time.
@@
Previous-token attention is measured the same as in the language model setting (Section˜C.2) with one difference: we compute the previous-token score not over a synthetic dataset but over a validation batch.
@@
Several attention heads are easy to identify by virtue of being both concentrated and consistent. These are depicted in Figure˜D.7 and include: (B1H3y) previous-token heads (also present in the language model case), (B1H1y) previous-x, and (B1H4x, B2H1y) previous-y heads. Other training runs also include self-attention heads.
[…]
@@
Suppose an attention head $h$ in block $b$ has the following constant attention pattern (after the softmax) $A^{(b,h)}=\sum_{i}\delta_{l(i)\,i}$. That is, for each token $i$, that attention head attends solely to a single earlier token $l(i)\leq i$ and no others. Restricting to single-head attention (the argument generalizes straightforwardly), the final contribution of this attention head to the residual stream is the following (Phuong & Hutter, 2022):
@@
[…]
@@ D.2.4 Unembedding collapse
where we order the outputs so that the $y$-token corresponds to the 0th row. Because we are reading out a single $y$ component, we can express the unembedding transformation in terms of “effective" unembedding weights and biases
@@
In Figure˜D.8, we plot $(\gamma,\beta)$, $((W_{U})_{[0,:]},(b_{U})_{[0]})$, and $(\tilde{W}_{U},\tilde{b}_{U})$ as a function of training steps, along with the mean weight over time. These are 64- and 1-dimensional vectors, so we can display the entire set of components. During stage LR3 the majority of weights $\beta$ and $W_{U}$ “collapse” to zero. Additionally, the layer normalization biases temporarily experience a large increase in variance before returning to small values. Despite this, the mean of the linear weights, layer normalization biases, and effective weights remains remarkably constant and close to zero throughout the entire process.
@@ D.2.5 Layer normalization collapse
The “collapse” in layer normalization weights is not unique to the unembedding. As depicted in Figure˜D.9, this behavior occurs in all layer norms except for the second MLP. The biases also remain centered close to zero even as the variance in biases grows much larger. Unlike in the unembedding, these layers begin to change earlier (starting halfway through LR2).
@@ D.2.6 Degeneracy and development
In the previous subsections, we provide a set of theoretical arguments for how embedding collapse (Section˜D.2.1), layer normalization collapse (Section˜D.2.5), and attention collapse (Section˜D.2.3) can lead to an increase in degeneracy, even while leaving the implemented function unchanged.
@@ Appendix E One-layer language model experiments
We also trained and ran some experiments on a one-layer language model (see Section˜F.1.1 for details). We aggregate results for the one-layer language model here, mirroring the experiments for the two-layer language model where possible. The early development of the one-layer model has many parallels with the two-layer model. At a single stage boundary, just as it occurs in the two-layer model, the one-layer model minimizes its bigram score (see Section˜C.1.1), begins utilizing the positional embedding to noticeably improve performance (see Section˜C.2.1), and starts making sudden improvements to the same $n$-gram scores (see Section˜C.1.2). Remarkably this occurs at the same checkpoint as in the 2-layer model (at 900 training steps).
@@
Figure E.1: We train a one-layer transformer model in the language setting to compare with the two-layer model. The development of certain behavioral and structural metrics over time closely mirrors the development of the same metrics in the early stages of the two-layer language model. Top: test loss and LLC estimations over time for the one-layer attention-only transformer, compare with Figure˜1(a). Bottom: bigram score, test loss with positional embedding ablated, and $n$-gram scores for the one-layer attention-only transformer, compare with Figure˜4(a,b,c).
Figure E.2: A more detailed version of Figure˜E.1 for the one-layer language model. Top: Loss, LLC, and weight norm, along with an overlaid Gaussian process fit to these curves (red dotted lines). Bottom: Associated slopes, both numerically estimated finite differences (transparent blue) and of the Gaussian process (red dotted lined).
@@ F.1.1 Architecture
The language model architectures we consider are one- and two-layer attention-only transformers. They have a context length of 1024, a residual stream dimension of $d_{model}=256$, $H=8$ attention heads per layer, and include layer normalization layers. We also used a learnable Shortformer positional embedding (Press et al., 2021). The resulting models have a total of $d=3,091,336$ parameters for $L=1$ and $d=3,355,016$ parameters for $L=2$. We used an implementation provided by TransformerLens (Nanda & Bloom, 2022).
@@
Token Embedding Weights 1,280,0001,280,0001 , 280 , 000
Positional Embedding Weights 262,144262,144262 , 144
Layer 1 Layer Norm Weights 256256256
Layer 1 Layer Norm Bias 256256256
Layer 1 Attention Query Weights 65,53665,53665 , 536
Layer 1 Attention Key Weights 65,53665,53665 , 536
Layer 1 Attention Value Weights 65,53665,53665 , 536
Layer 1 Attention Output Weights 65,53665,53665 , 536
Layer 1 Attention Query Bias 256256256
Layer 1 Attention Key Bias 256256256
Layer 1 Attention Value Bias 256256256
Layer 1 Attention Output Bias 256256256
Layer 2 Layer Norm Weights N/A 256256256
Layer 2 Layer Norm Bias N/A 256256256
Layer 2 Attention Query Weights N/A 65,53665,53665 , 536
Layer 2 Attention Key Weights N/A 65,53665,53665 , 536
Layer 2 Attention Value Weights N/A 65,53665,53665 , 536
Layer 2 Attention Output Weights N/A 65,53665,53665 , 536
Layer 2 Attention Query Bias N/A 256256256
Layer 2 Attention Key Bias N/A 256256256
Layer 2 Attention Value Bias N/A 256256256
Layer 2 Attention Output Bias N/A 256256256
Final Layer Norm Weights 256256256
Final Layer Norm Bias 256256256
Unembedding Weights 1,280,0001,280,0001 , 280 , 000
Unembedding Bias 5,0005,0005 , 000
@@ F.1.2 Tokenization
For tokenization, we used a truncated variant of the GPT-2 tokenizer that cut the original vocabulary of 50,000 tokens down to 5,000 (Eldan & Li, 2023) to reduce the size of the model. We think this may contribute to the prominence of the the plateau at the end of LM1: the frequency of bigram statistics depends on your choice of tokens, and a larger tokenizer leads to bigrams that are individually much less frequent.
@@ F.1.3 Training
The models are trained on a single epoch over 50,00050,00050 , 000 steps on $\sim$5 billion tokens using a resampled subset of the Pile (Gao et al., 2020; Xie et al., 2023) using a batch size of 100100100. A snapshot was saved every 101010 steps for a total of 500050005000 checkpoints, though a majority of analysis used checkpoints every 100 steps. The training time was around 6 GPU hours per model on an A100. Additional seeds were trained on v4 TPUs at around 1.5 TPU hours per model.
@@
$n$ Data # of training samples 5,000,0005,000,0005 , 000 , 000
$T$ Data # of training steps 50,00050,00050 , 000
@@
$L$ Model # of layers in the model 222
@@
$\eta$ Optimizer Learning rate 0.0010.0010.001
$\lambda_{\mathrm{wd}}$ Optimizer Weight Decay 0.050.050.05
$\beta_{1,2}$ Optimizer Betas (0.9,0.999)(0.9,0.999)( 0.9 , 0.999 )